6. The reaction, 2NO(g) + O2(g) → 2NO2(g), was studied and the following data were obtained.


[NO]0 [O2]0 Initial Rate
(molecules/cm3) (molecules/cm3) (molecules/cm3 s)
1.00 x1018 1.00 x1018 2.00 x1018
3.00 x1018 1.00 x1018 1.80 x1019
3.00 x1018 2.00 x1018 3.60 x1019

What would be the initial rate for an experiment where [NO]0 = 6.21 x 1018 molecules/cm3 and [O2]0 = 7.36 x 1018 molecules/cm3.
1
Expert's answer
2020-11-20T08:08:30-0500

To get the order of NO, hold O2 constant in experiment 1 and 2, you'll notice that as the reactant is tripled, the reaction increases by a factor of 9 which means it is 2nd order in NO

To get the order of O2, hold NO2 constant in experiment 2 and 3, you'll notice that as the reactants doubles, the reaction rate also doubles, which means it is 1st order in O2

Rate = k[NO]2[O2]

Let's take experiment 1 to find k

2.00 × 1018 /[ (1.0×1018 )2 (1.0×1018)] = K

K = 2.0 × 10-36


R = (2×10-36)×(6.21×1018 )2 (7.36 ×1018)

Rate= 5.68 ×1020 molecules/seconds cm3